# Point-hyperplane incidences via extremal graph theory

### v1

@@ 1.3 Large complete bipartite graphs in dense incidence graphs
On the other hand, in case $\varepsilon\ll\max\{m^{-\frac{1}{d}},n^{-\frac{1}{d-1}}\}$, Theorem 1.3 tells us that the largest complete bipartite subgraph might be just a star. Finally, observe that in case $m=n$, there is a large jump around $\epsilon\approx n^{-1/d}$, where $\operatorname{rs}_{d}(n,n,I)$ suddenly jumps from $\approx n^{1-1/d}$ to $\approx n^{1+1/d}$.
1.4 Forbidden induced patterns in incidence graphs
@@
The key idea used to prove our theoretical bounds is that incidence graphs of points and hyperplanes in $\mathbb{F}^{d}$ avoid a simple family $\mathcal{F}_{d}$ of graphs as induced subgraph. The family $\mathcal{F}_{d}$ includes all bipartite graphs on vertex classes $\{a_{1},\dots,a_{d}\}$ and $\{b_{1},\dots,b_{d}\}$, where $a_{i}b_{j}$ is an edge for $i\geq j-1$, and $a_{i}b_{i+2}$ is a non-edge for $i=1,\dots,d-2$.
@@
An alternative and perhaps easier way to think about the family is through the notion of a pattern. We define a pattern to be an edge labeling of a complete bipartite graph, in which every edge receives one of the labels 00, $1$ or ***. Every such pattern $\Pi$ defines a family of graphs $\mathcal{F}_{\Pi}$ on the same vertex set, where a graph $F$ is in the family if all edges of the pattern labeled by 00 do not appear in $F$, all edges labeled by $1$ appear in $F$, while the edges labeled by *** may or may not appear. To simplify the terminology, we say that a graph $G$ contains a pattern $\Pi$ if it contains an induced copy of some graph in the family $\mathcal{F}_{\Pi}$.
@@
The family $\mathcal{F}_{d}$ discussed above can be constructed in this way from the following pattern $\Pi_{d}$. Namely, $\Pi_{d}$ is a pattern on a balanced bipartite graph on vertices $a_{1},\dots,a_{d}$, $b_{1},\dots,b_{d}$ where $a_{i}b_{j}$ is labeled by $1$ for $i\geq j-1$ and $a_{i}b_{i+2}$ is labeled by 00, while all other edges are labeled by ***. See Figure 1 for an illustration.
@@
Assume, for the sake of contradiction, that $G=G(\mathcal{P},\mathcal{H})$ contains $\Pi_{d}$. Since the pattern $\Pi_{d}$ is symmetric, we may assume that the vertices $a_{1},\dots,a_{d}$ correspond to points $x_{1},\dots,x_{d}\in\mathcal{P}$, and the vertices $b_{1},\dots,b_{d}$ correspond to hyperplanes $H_{1},\dots,H_{d}\in\mathcal{H}$. We will prove by induction that $\dim\bigcap_{i=1}^{k}H_{i}=d-k$ for all $2\leq k\leq d$. This will suffice to derive the contradiction by observing that $x_{d-1},x_{d}\in\bigcap_{i=1}^{d}H_{i}$, which is at most a 00-dimensional space, i.e. a single point. This contradicts the assumption that $x_{1},\dots,x_{d}$ are distinct.
@@
For a linear algebraic perspective on the pattern $\Pi_{d}$, one might consider the $(d-2)\times(d-2)$ matrix whose rows and columns are indexed by $a_{1},\dots,a_{d-2}$ and $b_{3},\dots,b_{d}$, where the entry corresponding to $a_{i}$ and $b_{j}$ equals $1$ if the edge $a_{i}b_{j}$ is labeled 00 and equals 00 if the edge $a_{i}b_{j}$ is labeled $1$ (entries corresponding to the edges labeled by *** may be arbitrary). This defines an upper-triangular matrix with non-zero diagonal entries, a condition which ensures full rank. A similar family of graphs was also studied recently by Sudakov and Tomon [28] to establish Ramsey theoretic properties of algebraically defined graphs. Using the above lemma, we can just focus on the hereditary family of graphs that avoid the pattern $\Pi_{d}$. Then, we use the dependent random choice technique [16] and combinatorial methods to show that if $G$ is a $\Pi_{d}$-free graph with sufficiently many edges, then $G$ must contain a large complete bipartite graph.
In order to find configurations of points and hyperplanes matching our theoretical bounds, we consider optimal constructions of so called subspace evasive sets [10, 29]. We follow the ideas of [3, 5, 29] to transform large subspace evasive sets into point-hyperplane configuration with many incidences and no large bipartite subgraphs.
2 Illustration of our techniques in dimension 3
The purpose of this section is to introduce the main ideas of our paper in a relatively simple and concise way, while leaving the most general statements for later sections. We do this by focusing on the case $d=3$, and we begin by discussing the point-plane incidence bounds in case $G(\mathcal{P},\mathcal{H})$ does not contain $K_{s,s}$. We highlight that the forbidden pattern $\Pi_{3}$ is quite simple, since containing the pattern $\Pi_{3}$ is equivalent to containing the bipartite graph $K_{3,3}$ minus an edge as an induced subgraph.
Throughout this paper, we use standard graph theoretic notation. Given a graph $G$ and a vertex $v\in V(G)$, we denote by $\deg v$ the degree of $v$, and $N(v)$ the neighbourhood of $v$. Also, if $v_{1},\dots,v_{t}\in V(G)$, we denote by $N(v_{1},\dots,v_{t})$ the common neighbourhood of $v_{1},\dots,v_{t}$, i.e. $N(v_{1},\dots,v_{t})=N(v_{1})\cap\dots\cap N(v_{t})$.
@@ 2 Illustration of our techniques in dimension 3
In what follows, we give the proof of Theorem 1.3 in the case $d=3$, with an additional assumption that every point lies in many planes and every plane contains many points. This is mostly a technical assumption, which we will remove in the proof of the general case.
Lemma 2.2.
Let $\mathcal{P}$ be a set of $n$ points and $\mathcal{H}$ be a set of $m\geq n$ hyperplanes in $\mathbb{F}^{3}$. Assume that $I(\mathcal{P},\mathcal{H})\geq\varepsilon mn$ for some $\varepsilon$ satisfying $\varepsilon\geq 6\max\{n^{-1/2},m^{-1/3}\}$. Furthermore, assume that every point $x\in\mathcal{P}$ lies on at least $\varepsilon m/3$ hyperplanes of $\mathcal{H}$, and every hyperplane $H\in\mathcal{H}$ contains at least $\varepsilon n/3$ points of $\mathcal{P}$. Then $\operatorname{rs}(\mathcal{P},\mathcal{H})\geq\left(\frac{\varepsilon}{6}\right)^{2}mn$.
Proof.
Let $A$ and $B$ be the vertex classes of $G(\mathcal{P},\mathcal{H})$ with $|A|=n$ and $|B|=m$. Our proof relies on a simple observation that if $X=N(v_{1})\cap N(v_{2})$ for some two distinct vertices $v_{1},v_{2}\in A$, then any vertex $v_{3}\in A$ either satisfies $|N(v_{3})\cap X|\leq 1$ or $X\subseteq N(v_{3})$. Suppose this is not the case and there exists a vertex $v_{3}\in A$ with a non-neighbour $u_{1}\in X$ and two neighbours $u_{2},u_{3}\in X$. Then, the vertices $v_{1},v_{2},v_{3},u_{1},u_{2},u_{3}$ form the pattern $\Pi_{3}$, which is not possible due to Lemma 1.4.
Let us take vertices $v_{1},v_{2}\in A$ with the largest possible common neighbourhood $X=N(v_{1})\cap N(v_{2})$. By a similar computation as in the proof of Lemma 2.1 we get that the average size of the common neighbourhood of two vertices of $A$ is
$\frac{1}{\binom{n}{2}}\sum_{u\in B}\binom{\deg u}{2}\geq\frac{m}{\binom{n}{2}}\binom{\frac{1}{m}\sum_{u\in B}\deg u}{2}\geq\frac{\varepsilon^{2}m}{2},$ (2)
thus we have $|X|\geq\varepsilon^{2}m/2$.
Let $V_{1}$ be the set of vertices $v_{3}\in A$ for which $X\subseteq N(v_{3})$ and let $V_{2}$ be the set of vertices $v_{3}\in A$ for which $|N(v_{3})\cap X|\leq 1$. See Figure 2 for an illustration of these sets. By our previous observation, $A=V_{1}\cup V_{2}$. Let us now count the edges incident to $X$. Since every vertex of $X$ has degree at least $\varepsilon n/3$ by the conditions of our theorem, $X$ is incident to at least $\varepsilon n|X|/3$ edges. On the other hand, there are $|V_{1}||X|$ edges between $X$ and $V_{1}$, and at most $|V_{2}|$ edges between $X$ and $V_{2}$. Thus,
$\frac{\varepsilon}{3}n|X|\leq e(A,X)=e(V_{1},X)+e(V_{2},X)\leq|V_{1}||X|+|V_{2}|.$
Since $\varepsilon>6m^{-1/3}$, we have $\varepsilon n|X|/6\geq\varepsilon^{3}nm/12>n\geq|V_{2}|$, and so $|V_{1}||X|\geq\varepsilon n|X|/6$. Therefore, $|V_{1}|\geq\varepsilon n/6$. From this, we get the complete bipartite graph on the vertex set $V_{1}\cup X$ with $|X|\cdot|V_{1}|\geq\varepsilon^{3}mn/6$ edges. This is weaker than our promised bound, so let us improve it as follows.
As $X$ is the largest common neighbourhood of two vertices in $A$ and any two vertices of $V_{1}$ contain $X$ in their common neighbourhood, this shows that no two vertices of $V_{1}$ have a common neighbour outside $X$. In particular, every vertex outside $X$ sends at most $1$ edge to $V_{1}$, and so $e(V_{1},B)\leq|X||V_{1}|+m$. On the other hand, every vertex of $V_{1}$ has degree at least $\varepsilon m/3$, and therefore we have
$|V_{1}|\cdot\frac{\varepsilon}{3}m\leq e(V_{1},B)\leq|X||V_{1}|+m.$
Since $\varepsilon>6n^{-1/2}$, we have $|V_{1}|\cdot\varepsilon m/6\geq\frac{\varepsilon^{2}n}{36}m\geq m$ and so $|X||V_{1}|\geq|V_{1}|\cdot\varepsilon m/6$. Thus, we conclude that $|X|\geq\varepsilon m/6$ and so $G$ contains a complete bipartite subgraph between $V_{1}$ and $X$, with at least $|V_{1}||X|\geq\varepsilon^{2}mn/36$ edges. ∎
[…]
Figure 2: An illustration for the proof of Lemma 2.2
3 Point-hyperplane incidences with no $K_{s,s}$
In this section, we prove the upper bounds in Theorem 1.2. We prepare the proof with the following lemma about finding the pattern $\Pi_{d}$ under some simple conditions. We recall that $\Pi_{d}$ is the pattern on a balanced bipartite graph on vertices $a_{1},\dots,a_{d}$, $b_{1},\dots,b_{d}$ where $a_{i}b_{j}$ is labeled by $1$ for $i\geq j-1$ and $a_{i}b_{i+2}$ is labeled by 00, while all other edges are labeled by ***.

### v2

@@ 1.1 Point-hyperplane incidences
On the other hand, in case $\varepsilon\ll\max\{m^{-\frac{1}{d-1}},n^{-\frac{1}{d}}\}$, Theorem 1.3 tells us that the largest complete bipartite subgraph might be just a star. Finally, observe that in case $m=n$, there is a large jump around $\varepsilon\approx n^{-1/d}$, where $\operatorname{rs}_{d}(n,n,I)$ suddenly jumps from $\approx n^{1-1/d}$ to $\approx n^{1+1/d}$.
1.2 Point-variety incidences
The number of incidences between points and varieties in the real space has been extensively studied (see Section 4.1 for a gentle introduction on varieties). Pach and Sharir [39, 40] proved that if $\mathcal{P}$ is a set of $m$ points and $\mathcal{V}$ is a set of $n$ algebraic curves of degree at most $\Delta$ in $\mathbb{R}^{2}$ such that $G(\mathcal{P},\mathcal{V})$ is $K_{s,t}$-free, then
$I(\mathcal{P},\mathcal{V})=O_{s,t,\Delta}\left(m^{\frac{s}{2s-1}}n^{\frac{2s-2}{2s-1}}+m+n\right).$
Zahl [54] and Basu and Sombra [4] extended this bound to dimension 3 and 4, respectively. Subsequently, Fox, Pach, Sheffer, Suk, and Zahl [22] established the following general bound in $\mathbb{R}^{D}$: if $\mathcal{P}$ is a set of $m$ points and $\mathcal{V}$ is a set of $n$ varieties in $\mathbb{R}^{D}$ of degree at most $\Delta$, and $G(\mathcal{P},\mathcal{V})$ is $K_{s,t}$-free, then
$I(\mathcal{P},\mathcal{V})=O_{D,s,t,\Delta,\varepsilon}\left(m^{\frac{(D-1)s}{Ds-1}+\varepsilon}n^{\frac{D(s-1)}{Ds-1}}+m+n\right),$ (2)
for every $\varepsilon>0$. However, this bound is only known to be tight in the special case when $s=2$ and $t$ is sufficiently large, see [45].
Here, we consider a similar problem over arbitrary fields. That is, we consider the problem of bounding the number of incidences between points and varieties in $\mathbb{F}^{D}$, where $\mathbb{F}$ is an arbitrary field, under the assumption that the incidence graph is $K_{s,s}$-free (our proofs and results remain essentially the same if we consider the more general problem of forbidding $K_{s,t}$ for $t\geq s$, so we assume $s=t$ to simplify notation).
Theorem 1.4.
Let $\mathcal{P}$ be a set of $m$ points and let $\mathcal{V}$ be a set of $n$ varieties in $\mathbb{F}^{D}$, each of dimension $d$ and degree at most $\Delta$. If the incidence graph $G(\mathcal{P},\mathcal{V})$ is $K_{s,s}$-free, then
$I(\mathcal{P},\mathcal{V})\leq O_{d,\Delta,s}(m^{\frac{d}{d+1}}n+m).$
Let us compare our bound with (2). As $s\rightarrow\infty$, the bound in (2) approaches the shape $m^{\frac{D-1}{D}}n+m+n$, which matches the one in Theorem 1.4 in case $d=D-1$. For smaller $d$, Theorem 1.4 provides even better bounds. Curiously, our bound does not depend on the dimension $D$ of the ambient space, it only depends on the degree and the dimension of the varieties. Moreover, our bounds are essentially tight, as the next theorem shows.
Theorem 1.5.
Let $d<D$ be positive integers, $\alpha>0$, and $m,n$ be positive integers such that $n=\lfloor m^{\alpha}\rfloor$, and $m,n$ are sufficiently large with respect to $D$ and $\alpha$. Then there exist a prime $p$, a set of $m$ points $\mathcal{P}\subseteq\mathbb{F}_{p}^{D}$ and a set $\mathcal{V}$ of $n$ varieties in $\mathbb{F}_{p}^{D}$ of dimension $d$ and degree at most $\Delta=\lceil(1+\alpha)(d+1)\rceil^{2}$ such that $G(\mathcal{P},\mathcal{V})$ does not contain $K_{s,s}$ with $s=\Delta^{1/2}$, and
$I(\mathcal{P},\mathcal{V})\geq\Omega_{d,\alpha}(m^{\frac{d}{d+1}}n).$
We provide two proofs of Theorem 1.4. The first one uses the framework of induced Turán problems. We construct a bipartite graph $H_{d,\Delta}$ such that incidence graphs of points and varieties in $\mathbb{F}^{D}$ of dimension $d$ and degree at most $\Delta$ avoid $H_{d,\Delta}$ as an induced subgraph, and every vertex of $H_{d,\Delta}$ in one of the parts has degree at most $d+1$. Then, we adapt the ideas of [30] to prove bounds on the number of edges of induced $H_{d,\Delta}$-free and $K_{s,s}$-free bipartite graphs.
An alternative proof of Theorem 1.4 is obtained by establishing a bound on the VC-dimension of incidence graphs and the rate of growth of their shatter functions. To do this we bound the number of incidence-patterns between points and lower-dimensional varieties, extending the celebrated result of Rónyai, Babai and Ganapathy [42] on the number of zero-patterns of polynomials.
We remark that graphs with bounded VC-dimension avoid induced copies of a certain forbidden bipartite subgraph with similar properties as $H_{d,\Delta}$, but our explicit construction of $H_{d,\Delta}$ is more straightforward. Moreover, in certain situations, like the unit-distance problem for instance, finding specific forbidden bipartite graphs leads to better bounds compared to the ones which follow from VC-dimension arguments.
Finally, in order to prove Theorem 1.5 as well as other constructions in this paper, we use the results from [51], which are based on the random algebraic method originally introduced by [8].
1.3 Unit distances in finite fields
The celebrated Erdős unit distance problem [19, 20] is one of the most notorious open problems in combinatorial geometry, asking to estimate the function $f_{d}(n)$, the maximum number of unit distances spanned by $n$ points in $\mathbb{R}^{d}$. The cases $d=2,3$ are the most difficult ones, with a large gap remaining between the best known lower and upper bounds. For $d=2$, the state-of-the-art is $n^{1+\Omega(1/\log\log n)}<f_{2}(n)<O(n^{4/3})$, where the lower bound is due to Erdős [20], and the upper bound is due to Spencer, Szemerédi and Trotter [48]. For $d=3$, we know that $\Omega(n^{4/3}\log\log n)<f_{3}(n)<n^{1.498}$ by [20] and [55], respectively. On the other hand, for $d\geq 4$, it is not difficult to construct a set of points achieving $f_{d}(n)=\Theta_{d}(n^{2})$ [33]. Indeed, if $d\geq 4$, one can take two orthogonal linear subspaces $V_{1}$ and $V_{2}$, each of dimension at least 2. Then, placing $n/2$ points on the origin-centered sphere $S_{i}\subset V_{i}$ of radius $1/\sqrt{2}$ for $i=1,2$, we get at least $n^{2}/4$ unit distances. However, as it is shown in [22], in some sense this construction is the only way one can get a quadratic number of unit distances. More precisely, if we assume that no $s$ points are contained in a $(d-3)$-dimensional sphere, then the maximum number of unit distances is at most
$O_{d,s,\varepsilon}(n^{2-\frac{2}{d+1}+\varepsilon})$
for every $\varepsilon>0$ (see also [25] for further strengthening). The key observation is that such unit distance graphs are semi-algebraic of complexity at most $2$ containing no copy of $K_{s,s}$. The bounds on the number of edges of such graphs follow from the more general bounds (3) on Zarankiewicz’s problem for algebraic graphs, which will be discussed in the next section. In the context of finite fields, this problem also received attention when $d=3$. In this dimension, there are stronger natural restrictions on the unit-distance graphs, beyond being $K_{s,s}$-free. Under these conditions, Zahl [56] proved that a set of $n$ points determines at most $O(n^{3/2})$ unit distances assuming $-1$ is not a square in $\mathbb{F}$, while Rudnev [44] established the upper bound $O(n^{8/5})$ without this assumption.
Here, we completely resolve the problem of bounding the number of unit distances over arbitrary fields, assuming the unit distance graph is $K_{s,s}$-free. A unit sphere in $\mathbb{F}^{d}$ with center $a=(a_{1},\dots,a_{d})\in\mathbb{F}^{d}$ is defined as the set of points $(x_{1},\dots,x_{d})\in\mathbb{F}^{d}$ which satisfy
$(x_{1}-a_{1})^{2}+\dots+(x_{d}-a_{d})^{2}=1.$
Furthermore, points $a$ and $b$ are at unit distance if $b$ is contained in the unit sphere with center $a$.
Theorem 1.6.
Let $\mathcal{P}$ be a set of $n$ points in $\mathbb{F}^{d}$. If the unit distance graph of $\mathcal{P}$ does not contain $K_{s,s}$, then the number of unit distances spanned by $\mathcal{P}$ is at most $O_{d,s}(n^{2-\frac{1}{\lceil d/2\rceil+1}})$.
Furthermore, this bound is sharp in the following sense.
Theorem 1.7.
Let $n,d$ be positive integers. There exists a constant $s=s(d)$, a finite field $\mathbb{F}_{q}$ and a set of $n$ points $\mathcal{P}\subseteq\mathbb{F}_{q}^{d}$ such that the unit distance graph on $\mathcal{P}$ does not contain $K_{s,s}$ and $\mathcal{P}$ spans $\Omega_{d}(n^{2-\frac{1}{\lceil d/2\rceil+1}})$ unit distances.
1.4 Algebraic graphs
Zarankiewicz’s problem asks for the maximum number of edges in an $n$-vertex graph without a copy of $K_{s,s}$. Fox, Pach, Sheffer, Suk, and Zahl [22] considered a variant of this problem in which case the host graph is semi-algebraic. We say that a bipartite graph $G=(A,B,E)$ is semi-algebraic of description complexity $t$ in $(\mathbb{R}^{d_{1}},\mathbb{R}^{d_{2}})$, if $A\subset\mathbb{R}^{d_{1}}$, $B\subset\mathbb{R}^{d_{2}}$, and the edges are defined by the sign-patterns of $t$ polynomials $f_{1},\dots,f_{t}:\mathbb{R}^{d_{1}}\times\mathbb{R}^{d_{2}}\rightarrow\mathbb{R}$, each of degree at most $t$. In [22] it is proved that if such a graph $G$ is $K_{s,s}$-free with $|A|=m$ and $|B|=n$, then
$|E(G)|=O_{d_{1},d_{2},t,s,\varepsilon}\left(m^{\frac{d_{2}(d_{1}-1)}{d_{1}d_{2}-1}+\varepsilon}n^{\frac{d_{1}(d_{2}-1)}{d_{1}d_{2}-1}}+m+n\right)$
for every $\varepsilon>0$. In the special case of interest $d=d_{1}=d_{2}$, this gives
$|E(G)|=O_{d,t,s,\varepsilon}\left((mn)^{\frac{d}{d+1}+\varepsilon}+m+n\right).$ (3)
This bound is only known to be tight for $d=2$. It is worth highlighting that this upper bound matches (1) up to the $\varepsilon$ error term.
We consider the analogous problem for algebraic graphs, that is, graphs that are defined with respect to zero-patterns of polynomials over arbitrary fields. Given two sets of points $\mathcal{P}\subseteq\mathbb{F}^{d_{1}},\mathcal{Q}\subseteq\mathbb{F}^{d_{2}}$, a collection of $t$ polynomials $f_{1},\dots,f_{t}:\mathbb{F}^{d_{1}}\times\mathbb{F}^{d_{2}}\to\mathbb{F}$, and a boolean formula $\Phi$, we define the bipartite graph $G$ on the vertex set $\mathcal{P}\cup\mathcal{Q}$, where $x\in\mathcal{P}$ and $y\in\mathcal{Q}$ are adjacent if and only if
$\Phi([f_{1}(x,y)=0],\dots,[f_{t}(x,y)=0])=1.$
We say that $G$ is an algebraic graph of description complexity $t$ if it can be described by $t$ polynomials $f_{1},\dots,f_{t}$, each of which has degree at most $t$.
Theorem 1.8.
Let $G$ be an algebraic graph of description complexity at most $t$ on the vertex set $\mathcal{P}\cup\mathcal{Q}$, where $\mathcal{P}\subseteq\mathbb{F}^{d_{1}}$, $\mathcal{Q}\subseteq\mathbb{F}^{d_{2}}$ and $|\mathcal{P}|=m$, $|\mathcal{Q}|=n$. If $G$ is $K_{s,s}$-free, the number of edges in $G$ is at most
$O_{d_{1},d_{2},t,s}(\min\{m^{1-1/d_{1}}n,mn^{1-1/d_{2}}\}).$
Moreover, this bound is tight, for every $d_{1},d_{2}$, there exists $t$ such that if $m,n$ are sufficiently large and $s\geq d_{1}+d_{2}$, the following holds. There exist a field $\mathbb{F}$ and a $K_{s,s}$-free algebraic graph $G$ of description complexity at most $t$ in $(\mathbb{F}^{d_{1}},\mathbb{F}^{d_{2}})$ with parts of size $m,n$ and at least $\Omega(\min\{m^{1-1/d_{1}}n,mn^{1-1/d_{2}}\})$ edges.
Paper organization: In Section 2, we give a short illustration of our techniques discussing point-hyperplane incidences in 3-dimensional spaces, and introduce the key lemma connecting combinatorics and geometry. Then, in Section 3, we prove our general results on point-hyperplane incidences, Theorems 1.2 and Theorem 1.3, along with the constructions showing the tightness of these bounds. In Section 4 we discuss point-variety incidences and show two proofs of Theorem 1.4, while postponing the constructions until Section 6. The unit distance problem is discussed in Section 5, where the proofs of Theorems 1.6 and 1.7 are presented. In Section 6 we discuss the Zarankiewcz problem for algebraic graphs and show Theorems 1.5 and 1.8. We conclude our paper with a few remarks.
2 Forbidden induced patterns and illustration in dimension $3$
@@
The key idea used to prove our bounds is that incidence graphs of points and hyperplanes in $\mathbb{F}^{d}$ avoid a simple family $\mathcal{F}_{d}$ of graphs as induced subgraphs. The family $\mathcal{F}_{d}$ includes all bipartite graphs on vertex classes $\{a_{1},\dots,a_{d}\}$ and $\{b_{1},\dots,b_{d}\}$, where $a_{i}b_{j}$ is an edge for $i\geq j-1$, and $a_{i}b_{i+2}$ is a non-edge for $i=1,\dots,d-2$.
@@
An alternative and perhaps easier way to think about the family is through the notion of a pattern. We define a pattern to be an edge labeling of a complete bipartite graph, in which every edge receives one of the labels 00, $1$ or *. Every such pattern $\Pi$ defines a family of graphs $\mathcal{F}_{\Pi}$ on the same vertex set, where a graph $F$ is in the family if all edges of the pattern labeled by 00 do not appear in $F$, all edges labeled by $1$ appear in $F$, while the edges labeled by * may or may not appear. To simplify the terminology, we say that a graph $G$ contains a pattern $\Pi$ if it contains an induced copy of some graph in the family $\mathcal{F}_{\Pi}$.
@@
The family $\mathcal{F}_{d}$ discussed above can be constructed in this way from the following pattern $\Pi_{d}$. Namely, $\Pi_{d}$ is a pattern on a balanced bipartite graph on vertices $a_{1},\dots,a_{d}$, $b_{1},\dots,b_{d}$ where $a_{i}b_{j}$ is labeled by $1$ for $i\geq j-1$ and $a_{i}b_{i+2}$ is labeled by 00, while all other edges are labeled by *. See Figure 1 for an illustration.
@@
Assume, for the sake of contradiction, that $G=G(\mathcal{P},\mathcal{H})$ contains $\Pi_{d}$. Since the pattern $\Pi_{d}$ is symmetric, we may assume that the vertices $a_{1},\dots,a_{d}$ correspond to points $x_{1},\dots,x_{d}\in\mathcal{P}$, and the vertices $b_{1},\dots,b_{d}$ correspond to hyperplanes $H_{1},\dots,H_{d}\in\mathcal{H}$. We prove by induction that $\dim\bigcap_{i=1}^{k}H_{i}=d-k$ for all $2\leq k\leq d$. This suffices to derive the contradiction by observing that $x_{d-1},x_{d}\in\bigcap_{i=1}^{d}H_{i}$, which is at most a 00-dimensional space, i.e. a single point. This contradicts the assumption that $x_{1},\dots,x_{d}$ are distinct.
@@
For a linear algebraic perspective on the pattern $\Pi_{d}$, one might consider the $(d-2)\times(d-2)$ matrix whose rows and columns are indexed by $a_{1},\dots,a_{d-2}$ and $b_{3},\dots,b_{d}$, where the entry corresponding to $a_{i}$ and $b_{j}$ equals $1$ if the edge $a_{i}b_{j}$ is labeled 00 and equals 00 if the edge $a_{i}b_{j}$ is labeled $1$ (entries corresponding to the edges labeled by * may be arbitrary). This defines an upper-triangular matrix with non-zero diagonal entries, a condition which ensures full rank. A similar family of graphs was also studied recently by Sudakov and Tomon [50] to establish Ramsey theoretic properties of algebraically defined graphs. Using the above lemma, we can just focus on the hereditary family of graphs that avoid the pattern $\Pi_{d}$. Then, we use the dependent random choice technique [24] and combinatorial methods to show that if $G$ is a $\Pi_{d}$-free graph with sufficiently many edges, then $G$ must contain a large complete bipartite graph.
Let us now illustrate how Lemma 2.1 can be combined with the technique of dependent random choice to give a very short proof of Theorem 1.1 in dimension three. Note that when $d=3$, the forbidden pattern $\Pi_{3}$ is quite simple, since containing the pattern $\Pi_{3}$ is equivalent to containing the bipartite graph $K_{3,3}$ minus an edge as an induced subgraph.
Throughout this paper, we use standard graph theoretic notation. Given a graph $G$ and a vertex $v\in V(G)$, we denote by $\deg v$ the degree of $v$, and by $N(v)$ the neighbourhood of $v$. Also, if $v_{1},\dots,v_{t}\in V(G)$, we denote by $N(v_{1},\dots,v_{t})$ the common neighbourhood of $v_{1},\dots,v_{t}$, i.e. $N(v_{1},\dots,v_{t})=N(v_{1})\cap\dots\cap N(v_{t})$.
@@
3 Point-hyperplane incidences
3.1 Upper bounds for $K_{s,s}$-free incidence graphs
In this section, we prove the upper bounds in Theorem 1.2. We prepare the proof with the following lemma about finding the pattern $\Pi_{d}$ under some simple conditions. We recall that $\Pi_{d}$ is the pattern on a balanced bipartite graph on vertices $a_{1},\dots,a_{d}$, $b_{1},\dots,b_{d}$ where $a_{i}b_{j}$ is labeled by $1$ for $i\geq j-1$ and $a_{i}b_{i+2}$ is labeled by 00, while all other edges are labeled by *.